# spectrum and local weak convergence of sparse random uniform hypergraphs

### v1

@@ 2.4.1. Galton-Watson measure
The present work is partially inspired by the study of $r$-set random walks on $k$-uniform hypergraphs. For any fixed $r<k$, one may consider an $r$-set subsets of the vertex set as fundamental units or “roots”, rather than individual vertices. A random walk in this setting proceeds by transitioning from one $r$-set to another, provided they share a common hyperedge. This $r$-set perspective naturally arises in various contexts across the literature, particularly in the analysis of random walks, the construction of adjacency-type matrices, and spectral techniques based on hypergraph structure [45, 32, 41]. Formally, the $r$-set random walk on a $k$-uniform hypergraph may be interpreted as a random walk on the $r$-set-weighted line graph associated with the hypergraph. Notably, in the special case when $k=2$ and $r=1$, this construction reduces to the classical simple random walk on Erdős–Rényi random graphs. The Linial Meshulam complex [39] is a special case of this $r$-set perspective, where r=k1r=k-1.
@@ 2.5.1. Adjacency operators
If $f\in\ell_{c}(V(G))$, write $S_{f}=\operatorname{supp}(f)$; then $S_{f}$ is finite and $f(u)=0$ for all $u\notin S_{f}$. Hence for each $v\in V(G)$, $(A_{G}^{w}f)(v)$ can be nonzero only if there exists $u\in S_{f}$ with $w(v,u)\neq 0$. Therefore,
@@
is well defined, and since $\ell_{c}(V(G))$ is dense in $\ell^{2}(V(G))$, $A_{G}^{w}$ is densely defined on $\ell^{2}(V(G))$ with domain $\ell_{c}(V(G))$.
@@ 2.5.2. $r$-set adjacency operator
The $r$-set weighted line graph associated with a hypergraph $H$ is the weighted graph $G_{r}(H)=(\mathcal{V}_{r},\mathcal{E}_{r},w_{r,H})$, as defined in Section 2.1.4. We shall refer to the adjacency matrix of $G_{r}(H)$ as the $r$-set adjacency matrix of the hypergraph $H$. Suppose that $H$ is a locally finite $k$-uniform hypergraph. The space of finitely supported functions on $\mathcal{V}_{r}$ is denoted by $\ell_{c}(\mathcal{V}_{r})$, that is,
@@
The weight $w_{r,H}(S,S^{\prime})$ is defined in (1), and $w_{r,H}:\mathcal{V}_{r}\times\mathcal{V}_{r}\to\mathbb{R}$ is a weight function satisfying $w_{r,H}(S,S^{\prime})=w_{r,H}(S^{\prime},S)$ and $w_{r,H}(S,S^{\prime})=0$ if $S\nsim S^{\prime}$ (that is, there does not exists $e\in E(H)$ such that $S,S^{\prime}\subset e$). The weight function $w_{r,H}$ respect each $r$-set rooted hypergraph isomorphisms. That is, if $\sigma:V(H)\to V(H^{\prime})$ is an $r$-set rooted hypergraph isomorphism, then $w_{r,H}(S,S^{\prime})=w_{H^{\prime}_{r}}(\sigma(S),\sigma(S^{\prime}))$ for all $r$-cardinality subset $S,S^{\prime}$ of $V(H)$. For a finite $k$-uniform hypergraph $H$ on $n$ vertices, $A_{H}^{w_{r}}$ is represented by an $\binom{n}{r}\times\binom{n}{r}$ matrix. When $H\in H(n,k,p)$ is a random hypergraph, $A_{H}^{w_{r}}$ becomes a random matrix; we discuss its spectral properties further in this section.
@@ 2.6.2. Spectral measure of adjacency operator of weighted adjacency operator
For any graph $G$, the collection of square sumable function $\ell^{2}(V(G))$ is a Hilbert space with usual inner product $\langle f,g\rangle_{V(G)}=\sum_{v\in V(G)}f(v)g(v),$  $\mbox{ for }f,g\in\ell^{2}(V)$. That is, each $(G,o)\in\mathbb{G}^{*}_{w}$, corresponds to a Hilbert space $\ell^{2}(V(G))$. To incorporate the collection of Hilbert space $\{\ell^{2}(V(G)):(G,o)\in\mathbb{G}^{*}_{w}\}$ into a single Hilbert space we recall some notion from [3, Section-5, P.27]. Using any $\rho\in\mathcal{P}_{uni}(\mathbb{G}^{*}_{w})$ the global Hilbert space $\mathcal{H}$ is defined as the following a direct integral:
@@
Recall, $\nu_{d,\lambda}$ denotes the measure induced by the $d$-block Galton-Watson process with Poisson offspring distribution. Since the measure $\nu_{d,\lambda}$ belongs to $\mathcal{P}(\mathbb{G}^{*}_{w})$, the measure $\mu_{\nu_{d,\lambda}}$ is well defined. We show that under a suitable condition the limiting spectral measure of the $r$-adjacency matrices of $k$-uniform random hypergraphs is $\mu_{\nu_{d,\lambda}}$ with $d=\binom{k}{r}-1$. See 3.5.
@@ 3. Main results
Our first result is a natural extension of the local convergence of Erős-Rényi graphs in the $k$-uniform hypergraphs setup. Specifically, the result provides the 1-set local weak limit of $k$-uniform random hypergraphs. In this case, a root refers to a vertex of the hypergraphs..
@@
Let $\lambda>0$ and $G_{n}\in\mathcal{G}(n,p)$, the Erdős–Rényi graph. If $np\to\lambda$ as $n\to\infty$ then the local weak limit of $G_{n}$ is $\nu_{1,\lambda}$ almost surely converges. That is,
@@
The next result extends 3.1. It provides the $r$-set local weak convergence of $k$-uniform random hypergraphs for $1\leq r\leq k-1$.
@@
Though 3.1 follows directly from 3.3, we provide a separate proof of this result to enhance the readability of this work. For r=k1r=k-1, the (k1)(k-1)-set rooted $k$-uniform hypergraph resembles the Linial-Meshulam model [39, 42]. The notion of the $k$-set local limit of $H(n,k+1,p)$ matches with the notion of the local weak limit of the line graphs of the Linial-Meshulam model of dimension $k$, as defined in [2]. Thus, the following Corollary of 3.3 gives [2, Theorem 11].
@@
$\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}}\to\mu_{\nu_{d,\lambda}},\mbox{ as $n\to\infty$, almost surely}$
@@
In the last result, we assume that the expected degree of each vertex is $\lambda$, a fixed constant. In this paper, we do not attempt to find the LSD of $A_{H}^{w_{r}}$ when the expected degree of each vertex goes to infinity, that is, $\binom{n-r}{k-r}p=\lambda$ as $n\to\infty$. However, in this case, we believe that the limiting spectral distribution of $A_{H}^{w_{r}}$ is the semicircular law. This remains as our future work.
@@
Let $\lambda>0$ and $G_{n}\in\mathcal{G}(n,p)$, the Erdős-Rényi graphs, with $np\to\lambda$ as $n\to\infty$. The the LSD of the adjacency matrices of $G_{n}$ is $\Gamma_{1,\lambda}$} almost surely.
For $r=1$, the vertices are serving as roots. The adjacency matrix $A^{w_{1}}_{H}$ is generalization of some notion of conventional adjacency matrices associated with hypergraphs [9, 10]. An immediate Corollary of 3.5 is the following.
@@
In this section, we provide a proof of 3.1. Though this result directly follows from 3.3, since it is the simplest version of Benjamini-Schramm convergence in the case of $k$-uniform hypergraphs. We believe that proof will motivate the proof of the 3.3 and enhance the readability of the paper. The following two propositions are the key steps to prove the result.
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1 with $r=1$ and $H_{n}\in H(n,k,p)$. Then the sequence of expected measures of $\{U_{1}(H_{n})\}$ converges weakly to the measure $\nu_{k-1,\lambda}$. That is,
@@
Proof of 3.1.
The result follows from 4.1 and Section 4.2. ∎
@@ 4.1. Convergence in expectation
We present the proof of 4.1 in this section. Our proof is motivated by [16, Chapter 3]. First, we complete the proof using the following Proposition. For any weighted rooted graph $g$, the $\delta$-neighbourhood of the $1$-set rooted weighted-line graph is given by
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1 and $\mu_{n}=\mbox{\bf E}[U_{1}(H_{n})]$, where $H_{n}\in H(n,k,p)$. Then, for any $\delta>0$,
@@
where $\nu_{k-1,\lambda}$ be the induced measure of (k1)(k-1)-block Galton-Watson measure.
Proof of 4.1.
@@
Since $K$ is compact, for $\delta>0$, there exists finite collection of rooted graphs $\mathcal{S}(\subset\mathbb{G}^{*}_{w})$ such that $K\subseteq\bigcup\limits_{g\in\mathcal{S}}B_{\delta}(g)$. Thus, there exists $\delta>0$ such that
(10) $\sum\limits_{g\in\mathcal{S}}\nu_{k-1,\lambda}(B_{\delta}(g))\geq 1-\epsilon.$
Therefore, using 4.4, there exists $N_{o}\in\mathbb{N}$ such that
(11) $\sum\limits_{g\in\mathcal{S}}\mu_{n}(B_{\delta}(g))\geq 1-2\epsilon,\mbox{ for all }n\geq N_{0}.$
Let $f:\mathbb{G}_{w}^{*}\to\mathbb{R}$ is a uniformly continuous bounded function. Then
@@
$\leq$ $\sum\limits_{g\in\mathcal{S}}\|f\|_{\infty}|(\mu_{n}(B_{\delta}(g))-\nu_{k-1,\lambda}(B_{\delta}(g)))|+\int\limits_{\mathbb{G}_{w}^{*}\setminus\mathcal{S}}\|f\|_{\infty}(d\mu_{n}-d\nu_{k-1,\lambda}),$
where $\mathcal{S}=\bigcup\limits_{g\in\mathcal{S}}B_{\delta}(g)$. Observe that, by (10) and (11), we have
@@
Therefore, as $\|f\|_{\infty}$ is finite and $\epsilon$ is arbitrary, 4.4 implies that
@@
It remains to prove 4.4. The next subsection is dedicated for that.
4.1.1. Proof of 4.4
For the proof, we need to analyze the difference
(12) $\left|\mu_{n}(B_{\delta}((g))-\nu_{k-1,\lambda}(B_{\delta}(g)\right|.$
The following Lemma would help us to understand the meaning of $\mu_{n}(B_{\delta}((G_{1}(H_{n}),i)))$ for some $H_{n}\in H_{n}$.
@@
$\mu_{n}(A)$ $=\frac{1}{n}\sum\limits_{i\in[n]}\mbox{\bf E}(\delta_{(G_{1}(H_{n}(i)),,i)})(A)$
@@
where $G_{1}(GW_{k-1},o)$ denotes the 1-set line graph of $(GW_{k-1},o)_{\lceil\frac{1}{\epsilon}\rceil}$. Thus Equation 13 and (14) imply that the difference noted in (12) primarily arises from discrepancies in the neighbourhoods of $(G_{1}(H_{n})(1),1))$ and $(GW_{k-1},o)$. To analyze this more precisely, we now investigate the structure of the rooted random hypergraph $(H_{n},1)$, where $H_{n}\in H(n,k,p)$, which leads to the introduction of certain random variables that measure the extent of deviation between the respective neighbourhoods.

### v2

@@ 2.4.3. Structure of $(T_{k},S_{r})$
This work is motivated in part by the study of $r$-set random walks on $k$-uniform hypergraphs. For a $k$-uniform hypergraph $H$, the $r$-set random walk is the natural random walk on $G_{r}(H)$, where transitions occur between $r$-subsets that co-occur in a hyperedge of $H$. For any fixed $r<k$, one may consider $r$-element subsets of the vertex set as fundamental units or roots, rather than individual vertices. The walk transitions from $S_{1}$ to $S_{2}$ with positive probability if $S_{1}\cup S_{2}\subseteq e$ for some $e\in E(H)$; in other words, $S_{1}$ and $S_{2}$ are adjacent in $G_{r}(H)$. This $r$-set perspective appears in several contexts, particularly in the study of random walks, the construction of adjacency-type matrices, and spectral methods for hypergraphs [45, 32, 41]. Formally, the $r$-set random walk on a $k$-uniform hypergraph is equivalent to an weighted variation of random walk on its associated $r$-set-weighted line graph. Notably, in the special case when $k=2$ and $r=1$, this construction reduces to the classical simple random walk on Erdős–Rényi random graphs. The Linial–Meshulam complex [39] can also be viewed as a special case of this perspective, corresponding to r=k1r=k-1.
@@ 2.5.1. Adjacency operators
Here the weight is symmetric that is w(u,v)=w(v,u)w(u,v)=w(v,u)for any pair of vertices $u$ and $v$. If $f\in\ell_{c}(V(G))$, write $S_{f}=\operatorname{supp}(f)$; then $S_{f}$ is finite and $f(u)=0$ for all $u\notin S_{f}$. Hence for each $v\in V(G)$, $(A_{G}^{w}f)(v)$ can be nonzero only if there exists $u\in S_{f}$ with $w(v,u)\neq 0$. Therefore,
@@
is well defined, and since $\ell_{c}(V(G))$ is dense in $\ell^{2}(V(G))$, $A_{G}^{w}$ is densely defined on $\ell^{2}(V(G))$ with domain $\ell_{c}(V(G))$. Since $w$ is symmetric, the operator $A_{G}^{w}$ is symmetric on $\ell_{c}(V(G))$, hence densely defined and formally self-adjoint. If the vertex degrees are uniformly bounded and the weights are bounded, then $A_{G}^{w}$ extends to a bounded self-adjoint operator on $\ell^{2}(V(G))$.
@@ 2.5.2. $r$-set adjacency operator
The $r$-set weighted line graph associated with a hypergraph $H$ is the weighted graph $G_{r}(H)=(\mathcal{V}_{r},\mathcal{E}_{r},w_{r,H})$, as defined in Section 2.1.4. We shall refer to the adjacency matrix of $G_{r}(H)$ as the $r$-set adjacency matrix of the hypergraph $H$. As in the case of weighted graphs (see previous subsection), $A_{H}^{w_{r}}$ is initially defined on $\ell_{c}(\mathcal{V}_{r})$, making it a densely defined operator on $\ell^{2}(\mathcal{V}_{r})$. Suppose that $H$ is a locally finite $k$-uniform hypergraph. The space of finitely supported functions on $\mathcal{V}_{r}$ is denoted by $\ell_{c}(\mathcal{V}_{r})$, that is,
@@
Since the hypergraph $H$ is locally finite, each $r$-set $S$ participates in only finitely many hyperedges, so the set of $S^{\prime}$ with $w_{r,H}(S,S^{\prime})\neq 0$ is finite. This ensures the sums in (6) are finite. The weight $w_{r,H}(S,S^{\prime})$ is defined in (1), and $w_{r,H}:\mathcal{V}_{r}\times\mathcal{V}_{r}\to\mathbb{R}$ is a weight function satisfying $w_{r,H}(S,S^{\prime})=w_{r,H}(S^{\prime},S)$ and $w_{r,H}(S,S^{\prime})=0$ if $S\nsim S^{\prime}$ (that is, there does not exists $e\in E(H)$ such that $S,S^{\prime}\subset e$). The weight function $w_{r,H}$ respect each $r$-set rooted hypergraph isomorphism. That is, if $\sigma:V(H)\to V(H^{\prime})$ is an $r$-set rooted hypergraph isomorphism, then $w_{r,H}(S,S^{\prime})=w_{H^{\prime}_{r}}(\sigma(S),\sigma(S^{\prime}))$ for all $r$-cardinality subset $S,S^{\prime}$ of $V(H)$. For a finite $k$-uniform hypergraph $H$ on $n$ vertices, $A_{H}^{w_{r}}$ is represented by an $\binom{n}{r}\times\binom{n}{r}$ matrix. When $H\in H(n,k,p)$ is a random hypergraph, $A_{H}^{w_{r}}$ becomes a random matrix; we discuss its spectral properties further in this section.
@@ 2.6.2. Spectral measure of adjacency operator of weighted adjacency operator
For any graph $G$, the collection of square sumable function $\ell^{2}(V(G))$ is a Hilbert space with usual inner product $\langle f,g\rangle_{V(G)}=\sum_{v\in V(G)}f(v)\overline{g(v)},$  $\mbox{ for }f,g\in\ell^{2}(V)$. That is, each $(G,o)\in\mathbb{G}^{*}_{w}$, corresponds to a Hilbert space $\ell^{2}(V(G))$. To incorporate the collection of Hilbert space $\{\ell^{2}(V(G)):(G,o)\in\mathbb{G}^{*}_{w}\}$ into a single Hilbert space we recall some notion from [3, Section-5, P.27]. Using any $\rho\in\mathcal{P}_{uni}(\mathbb{G}^{*}_{w})$ the global Hilbert space $\mathcal{H}$ is defined as the following a direct integral:
@@
Recall, $\nu_{d,\lambda}$ denotes the measure induced by the $d$-block Galton-Watson process with Poisson offspring distribution. Since the measure $\nu_{d,\lambda}$ belongs to $\mathcal{P}(\mathbb{G}^{*}_{w})$, the measure $\mu_{\nu_{d,\lambda}}$ is well defined. We show that under a suitable condition the limiting spectral measure of the $r$-adjacency matrices of $k$-uniform random hypergraphs is $\mu_{\nu_{d,\lambda}}$ with $d=\binom{k}{r}-1$. See Theorem 3.5.
@@ 3. Main results
Our first result is a natural extension of the local convergence of Erős-Rényi graphs in the $k$-uniform hypergraphs setup. Specifically, the result provides the 1-set local weak limit of $k$-uniform random hypergraphs. In this case, a root refers to a vertex of the hypergraphs.
@@
Let $\lambda>0$ and $G_{n}\in\mathcal{G}(n,p)$, the Erdős–Rényi graph. If $np\to\lambda$ as $n\to\infty$ then the local weak limit of $G_{n}$ converges to $\nu_{1,\lambda}$ almost surely. That is,
@@
The next result extends Theorem 3.1. It provides the $r$-set local weak convergence of $k$-uniform random hypergraphs for $1\leq r\leq k-1$.
@@
Though Theorem 3.1 follows directly from Theorem 3.3, we provide a separate proof of this result to enhance the readability of this work. For r=k1r=k-1, the (k1)(k-1)-set rooted $k$-uniform hypergraph resembles the Linial-Meshulam model [39, 42]. The notion of the $k$-set local limit of $H(n,k+1,p)$ matches with the notion of the local weak limit of the line graphs of the Linial-Meshulam model of dimension $k$, as defined in [2]. Thus, the following Corollary of Theorem 3.3 gives [2, Theorem 11].
@@
$\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}}\rightsquigarrow\mu_{\nu_{d,\lambda}},\mbox{ as $n\to\infty$, almost surely}$
@@
In the last result, we assume that the expected degree of each vertex is $\lambda$, a fixed constant. In this paper, we do not attempt to find the LSD of $A^{w_{r}}_{H}$ when the expected degree of each vertex goes to infinity, that is, ${n-r\choose k-r}p=\lambda$ as $n\to\infty$. In this regime, it has been explored in [53] (for the case of 1-set adjacency matrices, with arguments that extend to general $r$-sets) that the limiting spectral distribution is semicircular, along with further results such as the convergence of edge eigenvalues. See also [56] for related developments in the sparse case.
@@
Let $\lambda>0$ and $G_{n}\in\mathcal{G}(n,p)$, the Erdős-Rényi graphs, with $np\to\lambda$ as $n\to\infty$. The LSD of the adjacency matrix of $G_{n}$ is $\Gamma_{1,\lambda}$} almost surely.
For $r=1$, the vertices are serving as roots. The adjacency matrix $A^{w_{1}}_{H}$ is generalization of some notion of conventional adjacency matrices associated with hypergraphs [9, 10]. An immediate Corollary of Theorem 3.5 is the following.
@@ 4. Convergence of $1$-set rooted uniform hypergraphs
In this section, we provide a proof of Theorem 3.1. Although the result follows directly from Theorem 3.3, it represents the simplest case of Benjamini-Schramm convergence for $k$-uniform hypergraphs. For this reason, we present an independent proof. We believe that proof will motivate the proof of the Theorem 3.3 and enhance the readability of the paper. The following two propositions are the key steps to prove the result.
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1 with $r=1$ and $H_{n}\in H(n,k,p)$. Then the sequence of expected measures ${\mbox{\bf E}}[U_{1}(H_{n})]$} converges weakly to the measure $\nu_{k-1,\lambda}$. That is,
@@
Proof of Theorem 3.1.
The result follows from Proposition 4.1 and Proposition 4.2. ∎
@@ 4.1. Convergence in expectation
We now prove Proposition 4.1, adapting ideas from [16, Chapter 3]. To this end, we begin with the following proposition. For any weighted rooted graph $g$, the $\delta$-neighbourhood of the $1$-set rooted weighted-line graph is given by
@@
Let $\lambda,k,r,n,p$ be as in Assumption 1 and $\mu_{n}=\mbox{\bf E}[U_{1}(H_{n})]$, where $H_{n}\in H(n,k,p)$. Thus, for any $\delta>0$,
@@
where $\nu_{k-1,\lambda}$ is the induced measure of (k1)(k-1)-block Galton-Watson measure.
Proof of Proposition 4.1.
@@
Since $K$ is compact, for $\delta>0$, there exists finite collection of rooted graphs ${S}(\subset\mathbb{G}^{*}_{w})$ such that $K\subseteq\bigcup\limits_{g\in{S}}B_{\delta}(g)$. Thus, there exists $\delta>0$ such that
(10) $\sum\limits_{g\in{S}}\nu_{k-1,\lambda}(B_{\delta}(g))\geq 1-\epsilon.$
Therefore, using Proposition 4.4, there exists $N_{o}\in\mathbb{N}$ such that
(11) $\sum\limits_{g\in{S}}\mu_{n}(B_{\delta}(g))\geq 1-2\epsilon,\mbox{ for all }n\geq N_{0}.$
Let $f:\mathbb{G}_{w}^{*}\to\mathbb{R}$ be a uniformly continuous bounded function. Then
@@
$\leq$ $\sum\limits_{g\in{S}}\|f\|_{\infty}|(\mu_{n}(B_{\delta}(g))-\nu_{k-1,\lambda}(B_{\delta}(g)))|+\int\limits_{\mathbb{G}_{w}^{*}\setminus\mathcal{S}}\|f\|_{\infty}(d\mu_{n}-d\nu_{k-1,\lambda}),$
where $\mathcal{S}=\bigcup\limits_{g\in{S}}B_{\delta}(g)$. Observe that, by (10) and (11), we have
@@
Therefore, as $\|f\|_{\infty}$ is finite and $\epsilon$ is arbitrary, Proposition 4.4 implies that
@@
It remains to prove Proposition 4.4. The next subsection is dedicated for that.
4.1.1. Proof of Proposition 4.4
For the proof, we need to analyze the deviation
(12) $\left|\mu_{n}(B_{\delta}(g))-\nu_{k-1,\lambda}(B_{\delta}(g))\right|.$
The following Lemma would help us to understand the meaning of $\mu_{n}(B_{\delta}((G_{1}(H_{n}),i)))$ for some $H_{n}\in H(n,k,p)$.
@@
$\mu_{n}(A)$ $=\frac{1}{n}\sum\limits_{i\in[n]}\mbox{\bf E}(\delta_{(G_{1}(H_{n}(i)),i)})(A)$
@@
where $G_{1}(GW_{k-1},o)$ denotes the 1-set line graph of $(GW_{k-1},o)_{\lceil\frac{1}{\epsilon}\rceil}$. Thus Equation (13) and (14) imply that the difference noted in (12) primarily arises from discrepancies in the neighbourhoods of $(G_{1}(H_{n})(1),1)$ and $(GW_{k-1},o)$. To analyze this more precisely, we now investigate the structure of the rooted random hypergraph $(H_{n},1)$, where $H_{n}\in H(n,k,p)$, which leads to the introduction of certain random variables that measure the extent of deviation between the respective neighbourhoods.